# Gradient correlation is needed to accelerate SGD with momentum

### v1

@@ Abstract
Empirically, it has been observed that adding momentum to Stochastic Gradient Descent (SGD) accelerates the convergence of the algorithm. However, the literature has been rather pessimistic, even in the case of convex functions, about the possibility of theoretically proving this observation. We investigate the possibility of obtaining accelerated convergence of the Stochastic Nesterov Accelerated Gradient (SNAG), a momentum-based version of SGD, when minimizing a sum of functions in a convex setting. We demonstrate that the average correlation between gradients allows to verify the strong growth condition, which is the key ingredient to obtain acceleration with SNAG. Numerical experiments, both in linear regression and deep neural network optimization, confirm in practice our theoretical results. Our code is available at https://github.com/Marien-RENAUD/Momentum_Stochastic_GD.
@@ 1 Introduction
Firstly, in the case of linear regression, depending on the data, SNAG can accelerate over SGD (Jain et al., 2018; Liu & Belkin, 2020; Varre & Flammarion, 2022). Unfortunately the methods used in those works are hardly generalizable outside of the linear regression case. On the other hand, in a convex setting, Vaswani et al. (2019) show that under an assumption over the gradient noise, named the Strong Growth Condition (SGC), the stochastic Nesterov algorithm can be stabilized, and it could accelerate over SGD. However, it is not clear which functions satisfy this assumption, and in which cases acceleration occurs.
@@
For many machine learning models, such as neural networks, the associated loss function is not convex (Li et al., 2018). However, even the question of the possibility to accelerate with SNAG in the convex setting is not solved yet. Finally, note that our core results (Propositions 1-2) do not assume convexity, and thus they could be used in non convex settings.
@@
(i) We complete a convergence result of Vaswani et al. (2019) with new almost sure convergence results (Theorem 4). (ii) We give a new characterization of the Strong Growth Condition (SGC) constant by using the correlation between gradients, quantified by RACOGA (Propositions 1-2), and we exploit this link to study the efficiency of SNAG. (iii) Using our framework, we study the theoretical impact of batch size on the algorithm performance, depending on the correlation between gradients (Theorem 5). (iv) We provide numerical experiments that show that RACOGA is a key ingredient to have good performances of SNAG compared to SGD.
@@ 2 Background
Assumption 2 implies that $f$ is $L$-smooth with $L\leq\frac{1}{N}\sum_{i=1}^{N}L_{i}\leq L_{max}$, and will be used for the convergence results for SGD (Theorems 1-2).
@@
The Nesterov accelerated gradient algorithm (Algorithm 7 in Appendix B.1) allows to achieve faster convergence than gradient descent when considering $L$-smooth functions that are convex or strongly convex, see Nesterov (1983; 2018). Intuitively, a momentum mechanism accelerates the gradient descent. A stochastic version of the Nesterov accelerated gradient algorithm can be developed, see Algorithm 2. Note that there exists several ways to write it (see Appendix B.2).
@@
To our knowledge, this assumption was introduced by Cevher & Vu (2019), as a relaxation of the maximal strong growth condition (Tseng, 1998; Solodov, 1998).
@@
The SGC implies interpolation (Assumption 1) if each $f_{i}$ is convex, see Remark 10 in Appendix G.3. However in the following results (Theorems 1-4), as we will only assume that the sum of $f_{i}$ is convex, SGC will not enforce interpolation. It will imply instead that minimizers of $f$ are critical points of all $f_{i}$.
@@ 3 Convergence speed of SNAG and comparison with SGD
In this section, we present convergence results for SNAG (Algorithm 2) under the Strong Growth Condition (SGC). Before doing so, we introduce convergence results for SGD (Algorithm 1) in order to compare the performance of these two algorithms.
@@ 3.1 Convergence results for SGD
Under Assumptions 1 and 2, SGD (Algorithm 1) guarantees the following asymptotic results:
@@
Note that in the strongly convex case, there is a need of averaging the trajectory along iterations. Proofs of Theorem 1 and Theorem 2 are in Appendix E.1.
@@ 3.2 Convergence in expectation for SNAG
Theorem 3 is a variation of Vaswani et al. (2019). Indeed, our proof (Appendix F) results in a slightly simpler formulation of the algorithm, as we do not have intermediate sequences of parameters. Note that we only use the $L$-smoothness of $f$ and the SGC instead of Assumptions 1-2 because SGC allows us to make weaker assumptions.
@@
In our comparison, we did not consider a convergence result for SGD (Algorithm 1) that assumes the SGC. We justify this choice in Appendix E.2.
@@ 3.3 Almost sure convergence for SNAG
We provide new asymptotic almost sure convergence results for SNAG (Algorithm 2).
@@
Assume $f$ is $L$-smooth , and that ~$\tilde{\nabla}_{K}$ verifies the SGC for $\rho_{K}\geq 1$. Then SNAG (Algorithm 2) guarantees the following asymptotic results:
@@
These bounds are asymptotically better than the finite time bounds, with $o(\frac{1}{n^{2}})$ compared to $\mathcal{O}(\frac{1}{n^{2}})$ (Theorem 3) for instance in the convex setting. A similar asymptotic speedup phenomenon happens in the deterministic setting (Attouch & Peypouquet, 2016).
@@ 4.1 Average positive correlation condition
Proposition 1 is proved in Appendix G.1. Without any assumption on $f$, Proposition 1 splits the norm of $\nabla f$ into two terms. One relies on the gradient estimator, while the other one involves the average correlation of gradients. From Proposition 1, we deduce the following consequence.
@@
In some cases, condition PosCorr could be too restrictive (see Appendix H). In the following section, we show how to ensure the SGC with a relaxed version of PosCorr, named RACOGA.
@@ 4.2 RACOGA: relaxing the PosCorr condition
Proposition 2 is proved in Appendix G.2. Note that RACOGA is always verified with $c=-\frac{1}{2}$, and we have $c\leq\frac{N-1}{2}$ (Appendix G.5).
@@
Proposition 2 creates a direct link between RACOGA and SGC. It indicates that the more correlation between gradients there is, the lower is the SGC constant, which results in improved convergence bounds for SNAG, see Theorems 3-4.
@@ 4.3 The strong growth condition with batch size $1$ determines how the performance scales with batch size
Lemma 1 is proved in Appendix G.3. It shows that if the SGC is verified for batches of size $1$, it is verified for any size of batch $K$ and we can compute an estimation of $\rho_{K}$.
@@
Compared to the theorems of section 3, Theorem 5 gives a bound on the number of $\nabla f_{i}$ we have to evaluate, not the number of iterations of the algorithm, see the proof in Appendix G.4.
@@ 5 Numerical experiments
Our theory indicates that the correlation between gradients, evaluated through RACOGA is needed to have SNAG (Algorithm 2) outperforming SGD (Algorithm 1). We provide numerical experiments to validate this statement by running SNAG to optimize classic machine learning models such as linear regression (Section 5.1) or classification neural network (Section 5.2). We also compare its performance with its deterministic version NAG (Algorithm 7), together with GD (Algorithm 6).
@@ 5.1 Linear regression
We consider the overparameterized case, i.e. $d>N$. Note that the linear regression problem is convex and smooth, so we are in the theoretical setting of this paper. Moreover, we can directly compute the parameters involved in the algorithms, see details in Appendix A.1.
@@
(a) No correlation within data
(b) Correlation within data
Figure 1: Illustration of the convergence speed of GD (Algorithm 6), SGD (Algorithm 1, batch size $1$), NAG (Algorithm 7) and SNAG (Algorithm 2, batch size $1$) on a linear regression problem, together with an histogram distribution of RACOGA values along the iterations of SNAG. Stochastic algorithms results are averaged over ten runs. On the left, data are generated by a law that make them fewly correlated, while on the right the data are generated by a gaussian mixture, leading to higher correlation. Note that the data correlation results in better performance of SNAG, whereas uncorrelated data lead to smaller RACOGA values, reducing the benefit of using SNAG.
On Figure 1(a), we observe that the lack of correlation inside data leads to very small values of RACOGA. It coincides with SGD being faster than SNAG. On the other hand, on Figure 1(b), we see that the presence of correlation inside the data makes the optimization path crosses areas with higher RACOGA values, allowing SNAG to be faster than SGD. These experimental results support our theoretical findings of Sections 3.
@@ 5.2 Neural networks
We use the SNAG version implemented in Pytorch (Algorithm 8) that is equivalent to Algorithm 2, see Appendix B.2
@@
Figure 2: Illustration of the convergence speed of GD (Algorithm 6), SGD (Algorithm 1, batch size $64$), NAG (Algorithm 8) and SNAG (Algorithm 3, batch size $64$) averaged over $10$ different initializations, together with an histogram distribution of RACOGA values taken along the optimization path, averaged over $10$ different initialisations, where the $x$-axis scale is logarithmic. On the left, we use a MLP to classify data sampled from a law such that they are fewly correlated. On the right we use a CNN to classify CIFAR10 images. Note that contrarily to Figure 1, the presence of correlation within data no longer influence the RACOGA values, that remains high in both cases, resulting in better performances of SNAG.
@@ Acknowledgments
This work was supported by the ANR project PEPR PDE-AI, and the French Direction Générale de l’Armement. Experiments presented in this paper were carried out using the PlaFRIM experimental testbed, supported by Inria, CNRS (LABRI and IMB), Universite de Bordeaux, Bordeaux INP and Conseil Regional d’Aquitaine (see https://www.plafrim.fr). We thank Bilel Bensaid for bringing to our knowledge the works about gradient diversity and gradient confusion, and Eve Descomps for her help for figure design.
@@
These supplements contain additional details on the numerical experiments, proofs of our theoretical results, and some additional insights. In Supplement A, we present additional details about numerical experiments and supplementary experiments. In Supplement B, an optimization background is provided for completeness. In Supplement C, we present and comment related works. In Supplement D, we provide simple examples of functions that do not verify SGC, or verify it for a large constant. In Supplement E, we provide convergence proofs for SGD (Algorithm 1). In Supplement F, we present convergence proof for SNAG (Algorithm 2). In Supplement G, we provide the proof of the results presented in Section 4. Finally in Supplement H, we give deeper explanations concerning the link between RACOGA and algorithms considered in this paper when applied to the problem of linear regression.

### v2

# Gradient correlation is a key ingredient to accelerate SGD with momentum

@@ Abstract
Empirically, it has been observed that adding momentum to Stochastic Gradient Descent (SGD) accelerates the convergence of the algorithm. However, the literature has been rather pessimistic, even in the case of convex functions, about the possibility of theoretically proving this observation. We investigate the possibility of obtaining accelerated convergence of the Stochastic Nesterov Accelerated Gradient (SNAG), a momentum-based version of SGD, when minimizing a sum of functions in a convex setting. We demonstrate that the average correlation between gradients allows to verify the strong growth condition, which is the key ingredient to obtain acceleration with SNAG. Numerical experiments, both in linear regression and deep neural network optimization, confirm in practice our theoretical results.
@@ 1 Introduction
Firstly, in the case of linear regression, depending on the data, SNAG can accelerate over SGD (Jain et al., 2018; Liu & Belkin, 2020; Varre & Flammarion, 2022). Unfortunately the methods used in those works are hardly generalizable outside of the linear regression case. On the other hand, in a convex setting, Vaswani et al. (2019) show that under an assumption over the gradient noise, named the Strong Growth Condition (SGC), SNAG can be stabilized, and it could accelerate over SGD. However, it is not clear which functions satisfy this assumption, and in which cases acceleration occurs.
@@
For many machine learning models, such as neural networks, the associated loss function is not convex (Li et al., 2018).
However, even in the convex setting, there is still work to do concerning the possibility of accelerating SGD with SNAG. For example, up to our knowledge, characterizing convex smooth functions that satisfy SGC has not been addressed yet.
Finally, note that our core results about gradient correlation (Propositions 1-2) do not assume convexity, and thus could be used in future works beyond the convex setting.
@@
(i) We give a new characterization of the Strong Growth Condition (SGC) constant by using the correlation between gradients, quantified by RACOGA (Propositions 1-2), and we exploit this link to study the efficiency of SNAG. (ii) Using our framework, we study the theoretical impact of batch size on the algorithm performance, depending on the correlation between gradients (Theorem 5). (iii) We complete convergence results of Vaswani et al. (2019); Gupta et al. (2023) with new almost sure convergence rates (Theorem 4). (iv) We provide numerical experiments that show that RACOGA is a key ingredient to have good performances of SNAG compared to SGD.
@@ 2 Background
Assumption 2 implies that $f$ is $L$-smooth with $L\leq\frac{1}{N}\sum_{i=1}^{N}L_{i}\leq L_{(K)}\leq L_{max}$, and will be used for the convergence results for SGD (Theorems 1-2).
@@
The Nesterov accelerated gradient algorithm (Algorithm 7 in Appendix B.1) allows to achieve faster convergence than gradient descent when considering $L$-smooth functions that are convex or strongly convex, see Nesterov (1983; 2018). Intuitively, a momentum mechanism accelerates the gradient descent. As proposed in Nesterov (2012), a stochastic version of the Nesterov accelerated gradient algorithm can be developed, see Algorithm 2. Note that there exists several ways to write it (see Appendix B.2).
@@
To our knowledge, this assumption was introduced in Polyak (1987), and further used by Cevher & Vu (2019) as a relaxation of the maximal strong growth condition (Tseng, 1998; Solodov, 1998).
@@
The SGC implies interpolation (Assumption 1) if each $f_{i}$ is convex, see Remark 11 in Appendix H.3. However in the following results (Theorems 1-4), as we will only assume that the sum of $f_{i}$ is convex, SGC will not enforce interpolation. It will imply instead that minimizers of $f$ are critical points of all $f_{i}$.
@@ 3 Convergence speed of SNAG and comparison with SGD
In this section, we present convergence results for SNAG (Algorithm 2) under the Strong Growth Condition (SGC). Before doing so, we introduce convergence results for SGD (Algorithm 1) in order to compare the performance of these two algorithms. All the results introduced in this section are summarized in Table 1.
@@
We denote by $\Omega$ the set of realization of the noise. We say that an algorithm $\{x_{n}\}_{n}$ converges almost surely with a rate negligible compared to $a_{n}\in{\mathbb{R}}_{++}^{{\mathbb{N}}}$, denoted by $f(x_{n})-f^{\ast}\overset{a.s.}{=}o\left(a_{n}\right)$, if and only if $\exists A\subset\Omega$, such that $\mathbb{P}(A)=1$ and $\forall\omega\in A$, $\forall\epsilon>0$, $\exists n_{0}\in{\mathbb{N}}$, such that $\forall n\geq n_{0}$,
$|f(x_{n}(\omega))-f^{\ast}|\leq\epsilon a_{n}.$ (3)
Algorithm Assumption over $f$ Convergence
SGD convex $\mathbb{E}$, $\varepsilon$ solution $\mathcal{O}\left(\frac{L_{(K)}}{\varepsilon}\right)$, Thm 1
strongly convex $\mathbb{E}$, $\varepsilon$ solution $\mathcal{O}\left(\frac{L_{(K)}}{\mu}\log\left(\frac{1}{\varepsilon}\right)\right)$, Thm 1
convex a.s., c.r $o\left(\frac{1}{n}\right)$, Thm 2
strongly convex a.s., c.r $o\left((1-\frac{\mu}{L}+\varepsilon^{\prime})^{n}\right)$, Thm 2
SNAG Convex $\mathbb{E}$, $\varepsilon$ solution $\mathcal{O}\left(\rho_{K}\sqrt{\frac{L}{\varepsilon}}\right)$, Thm 3
Strongly Convex $\mathbb{E}$, $\varepsilon$ solution $\mathcal{O}\left(\rho_{K}\sqrt{\frac{L}{\mu}}\log\left(\frac{1}{\varepsilon}\right)\right)$, Thm 3
Convex a.s., c.r $o\left(\frac{1}{n^{2}}\right)$, Thm 4
Strongly a.s., c.r $o\left((1-\frac{1}{\rho_{K}}\sqrt{\frac{\mu}{L}}+\varepsilon^{\prime})^{n}\right)$, Thm 4
Table 1: Summary of all the convergence results presented in Section 3. Results stated as $\epsilon$-solution refer to convergence results of the form of Equation (2). $c.r.$ stands for convergence rate. These results are stated as an upper bound of the form $f(x_{n})-f^{\ast}=\mathcal{O}\left(\psi_{n}\right)$, where $\psi_{n}$ is a sequence decreasing to 00.
@@ 3.1 Convergence results for SGD
Under Assumptions 1 and 2, SGD (Algorithm 1) guarantees, in the sense of (3), the following asymptotic results, :
@@
Note that in the convex case, there is a need of averaging the trajectory along iterations. Proofs of Theorem 1 and Theorem 2 are in Appendix E.1.
@@ 3.2 Convergence in expectation for SNAG
Theorem 3 is a variation of Vaswani et al. (2019); Gupta et al. (2023), see Appendix C.3. Indeed, our proof (Appendix G) leads to the same convergence result as Vaswani et al. (2019), although resulting in a slightly simpler formulation of the algorithm, as we do not have intermediate sequences of parameters, see Appendix C. Note that we only use the $L$-smoothness of $f$ and the SGC instead of Assumptions 1-2 because SGC allows us to make weaker assumptions.
@@
In our comparison, we neither considered a convergence result for SGD (Algorithm 1) that assumes the SGC, nor considered a result for SNAG (Algorithm 2) that does not assume SGC. In both cases, doing so would lead to misleading comparisons, see summary in Remark 9.
@@ 3.3 Almost sure convergence for SNAG
We provide new asymptotic almost sure convergence results for SNAG (Algorithm 2). Almost sure convergence has already been addressed in Gupta et al. (2023) without convergence rates.
@@
Assume $f$ is $L$-smooth , and that ~$\tilde{\nabla}_{K}$ verifies the SGC for $\rho_{K}\geq 1$. Then SNAG (Algorithm 2) guarantees, in the sense of (3), the following asymptotic results:
@@
See proofs in Appendix G. These bounds are asymptotically better than the finite time bounds, with $o(\frac{1}{n^{2}})$ compared to $\mathcal{O}(\frac{1}{n^{2}})$ (Theorem 3) for instance in the convex setting. A similar asymptotic speedup phenomenon happens in the deterministic setting (Attouch & Peypouquet, 2016).
@@ 4.1 Average positive correlation condition
Proposition 1 is proved in Appendix H.1. Without any assumption on $f$, Proposition 1 splits the norm of $\nabla f$ into two terms. One relies on the gradient estimator, while the other one involves the average correlation of gradients. From Proposition 1, we deduce the following consequence.
@@
In some cases, condition PosCorr could be too restrictive (see Appendix I). In the following section, we show how to ensure the SGC with a relaxed version of PosCorr, named RACOGA.
@@ 4.2 RACOGA: relaxing the PosCorr condition
Proposition 2 is proved in Appendix H.2. Note that RACOGA is always verified with $c=-\frac{1}{2}$, and we have $c\leq\frac{N-1}{2}$ (Appendix H.5).
@@
Proposition 2 creates a direct link between RACOGA and SGC. It indicates that the more correlation between gradients there is, the lower is the SGC constant, which results in improved convergence bounds for SNAG, see Theorems 3-4. Importantly if the gradients are too anti correlated, Proposition 2 could only be verified with $c$ arbitrary close to $-\frac{1}{2}$, resulting in a bound for $\rho_{1}$ increasing to +$\infty$+ ∞.
@@ 4.3 The strong growth condition with batch size $1$ determines how the performance scales with batch size
Lemma 1 is proved in Appendix H.3. It shows that if the SGC is verified for batches of size $1$, it is verified for any size of batch $K$ and we can compute an estimation of $\rho_{K}$.
@@
Compared to the theorems of section 3, Theorem 5 gives a bound on the number of $\nabla f_{i}$ we have to evaluate, not the number of iterations of the algorithm, see the proof in Appendix H.4.
@@ 5 Numerical experiments
Our theory indicates that the correlation between gradients, evaluated through RACOGA is needed to have SNAG (Algorithm 2) outperforming SGD (Algorithm 1). We provide numerical experiments to validate this statement by running SNAG to optimize classic machine learning models such as linear regression (Section 5.1) or classification neural network (Section 5.2). We also compare its performance with its deterministic version NAG (Algorithm 7), together with GD (Algorithm 6). Our code is available at https://github.com/J-Hermant/Momentum_Stochastic_GD.
@@ 5.1 Linear regression
We consider the overparameterized case, i.e. $d>N$. Note that the linear regression problem is convex and smooth, so we are in the theoretical setting of this paper. Moreover, we can directly compute the parameters involved in the algorithms except for a parameter $\lambda$ that replaces the unknown $\rho_{K}$ constant in the case of SNAG, see details in Appendix A.1.
@@
(a) Low correlation within data
(b) High correlation within data
Figure 1: Illustration of the convergence speed of GD (Algorithm 6), SGD (Algorithm 1, batch size $1$), NAG (Algorithm 7) and SNAG (Algorithm 2, batch size $1$) on a linear regression problem, together with an histogram distribution of RACOGA values along the iterations of SNAG. Stochastic algorithms results are averaged over ten runs. On the left, data are generated by a law that make them fewly correlated, while on the right the data are generated by a gaussian mixture, leading to higher correlation. The $\lambda$ parameter replaces the unknown SGC constant in the algorithm, see Appendix A.1. Note that the data correlation results in better performance of SNAG, whereas uncorrelated data lead to smaller RACOGA values, reducing the benefit of using SNAG.
SGD vs SNAG
Comparing right parts of Figure 1(a) and Figure 1(b), we observe that the lack of correlation inside data leads to smaller values of RACOGA. In the case of Figure 1(a), these lower RACOGA values coincide with SGD being faster than SNAG. On the other hand we see on Figure 1(b) that the presence of correlation inside the data makes the optimization path crosses areas with higher RACOGA values, allowing SNAG to be faster than SGD. These experimental results support our theoretical findings.
Deterministic vs stochastic
@@
Role of $\lambda$
In the case of SNAG (Algorithm 2), the choice of parameters from Theorem 3 involves the $\rho_{K}$ constant, that we do not know. We thus replace $\rho_{K}$ by a parameter $\lambda$. The higher the RACOGA are, the smaller $\rho_{K}$ is, and the smaller $\lambda$ can be chosen which results to more aggressive steps of the algorithm, as $s=\frac{1}{L\lambda}$. See details in Appendix A.1.
@@ 5.2 Neural networks
We use the SNAG version implemented in Pytorch (Algorithm 3), that is equivalent to Algorithm 2, see Appendix B.2
@@
Figure 2: Illustration of the convergence speed of GD (Algorithm 6), SGD (Algorithm 1, batch size $64$), NAG (Algorithm 3, full batch) and SNAG (Algorithm 3, batch size $64$) averaged over $10$ different initializations, together with an histogram distribution of RACOGA values taken along the optimization path, averaged over $10$ different initialisations, where the $x$-axis scale is logarithmic. On the left, we use a MLP to classify data sampled from a law such that they are fewly correlated. On the right we use a CNN to classify CIFAR10 images. Note that contrarily to Figure 1, the presence of correlation within data no longer influence the RACOGA values, that remains high in both cases, resulting in better performances of SNAG.
@@ Acknowledgments
This work was supported by the ANR project PEPR PDE-AI, and the French Direction Générale de l’Armement. Experiments presented in this paper were carried out using the PlaFRIM experimental testbed, supported by Inria, CNRS (LABRI and IMB), Universite de Bordeaux, Bordeaux INP and Conseil Regional d’Aquitaine (see https://www.plafrim.fr). We thank Bilel Bensaid to bringing to our knowlegde the works about gradient diversity and gradient confusion, and Eve Descomps for her help for figure design.
@@
These supplements contain additional details on the numerical experiments, proofs of our theoretical results, and some additional insights. In Supplement A, we present additional details about numerical experiments and supplementary experiments. In Supplement B, an optimization background is provided for completeness. In Supplement C, we present and comment related works. In Supplement D, we provide simple examples of functions that do not verify SGC, or verify it for a large constant. In Supplement E, we provide convergence proofs for SGD (Algorithm 1). In Supplement F, we introduce new convergence results for SNAG (Algorithm 2) without assuming that the SGC holds. In Supplement G, we present convergence proof for SNAG (Algorithm 2). In Supplement H, we provide the proof of the results presented in Section 4. Finally in Supplement I, we give deeper explanations concerning the link between RACOGA and algorithms considered in this paper when applied to the problem of linear regression.
@@ A.1 Linear regression
Under-parameterization with a realistic dataset
The Boston dataset (Harrison & Rubinfeld, 1978) concerns house pricing in the Boston area. This dataset can be used for linear regression, where we aim to predict house prices using $13$ variables such as tax rates. There are $200$ data, so $N=200$ and $d=13$, which means we are in an under-parameterized regime. In this case, we can hardly expect SGC to hold, as the noise of the gradient has no reason to be zero at the minimizers of $f$. On Figure 4, we tackle the linear regression problem on this dataset, and we observe that despite the under-parameterized setting, SNAG (Algorithm 2) appears to outperform GD (Algorithm 6), NAG (Algorithm 7) and SGD (Algorithm 1). RACOGA values can be either very large (e.g. for aggressive, i.e. small $\lambda$), but also close to $-\frac{1}{2}$, which corresponds to high anti-correlation. In this under-parameterized regime, the role of RACOGA is less clear. A possible explanation of the good performance of SNAG could be the variance reduction effect of momentum, see Gao et al. (2024).
Figure 4: Illustration of the convergence speed of GD (Algorithm 6), NAG (Algorithm 7), SGD (Algorithm 1) and SNAG (Algorithm 2) for the linear regression problem with the Boston dataset. The parameters of SNAG are choosen accordingly to Equation 16. Some RACOGA values are high, while other are also close to the worst anti-correlation case $-\frac{1}{2}$. Note that SNAG appears to converge faster than the other algorithms when choosing $\lambda$ small enough.